22.0-mL sample of 0.122 M diprotic acid (H2A) solution is titrated with 0.1016 M KOH. The acid ionization constants for the acid are Ka1=5.2×10−5 and Ka2=3.4×10−10.

A. At what added volume of base does the first equivalence point occur? ____ mL

B. At what added volume of base does the second equivalence point occur? _____ mL

12
Answer #1

A)
Balanced chemical equation is:
H2A + KOH ---> KHA + H2O


Here:
M(H2A)=0.122 M
M(KOH)=0.1016 M
V(H2A)=22.0 mL

According to balanced reaction:
1*number of mol of H2A =1*number of mol of KOH
1*M(H2A)*V(H2A) =1*M(KOH)*V(KOH)
1*0.122 M *22.0 mL = 1*0.1016M *V(KOH)
V(KOH) = 26.42 mL

Answer: 26.42 mL

B)
Balanced chemical equation is:
H2A + 2 KOH ---> K2A + 2 H2O


Here:
M(H2A)=0.122 M
M(KOH)=0.1016 M
V(H2A)=22.0 mL

According to balanced reaction:
2*number of mol of H2A =1*number of mol of KOH
2*M(H2A)*V(H2A) =1*M(KOH)*V(KOH)
2*0.122 M *22.0 mL = 1*0.1016M *V(KOH)
V(KOH) = 52.83 mL

Answer: 52.83 mL